Your age is now 7 years.your brother's age is 17 years now .after how many year your brother's age will 2 times of your age?
Answer by rothauserc(4718)

let x be the number of years when your brother is twice your age
2(7 + x) = 17 + x
14 + 2x = 17 + x
x = 3
In three years you will be 10 and your brother will be 20